OpenStudy (anonymous):

Tell whether the lines for each pair of equations is parallel, perpandicular, or neither. 1. Y=-3x+7 and -2x+6y=3 2. Y= -1/4x+10 and -2x+8y=6

OpenStudy (anonymous):

y = -3x + 7 -- slope here is -3 -2x + 6y = 3 (change to slope intercept form) 6y = 2x + 3 y = 2/6x + 3/6 y = 1/3x + 1/2 -- slope here is 1/3 These lines are perpendicular because the lines have negative reciprocal slopes. To find the negative reciprocal of a slope, just " flip " the slope and change the sign. The negative reciprocal of -3 is 1/3 (you see how I flipped the slope and changed the sign). ==================== y = -1/4x + 10 -- slope here is -1/4 -2x + 8y = 6 (change to slope intercept form) 8y = 2x + 6 y = 2/8x + 6/8 y = 1/4x + 3/4 -- slope here is 1/4 These lines are neither perpendicular or parallel. Parallel lines will have the same slope and perpendicular lines will have negative reciprocal slopes.